OpenStudy (anonymous):

The population of a type of local dragonfly can be found using an infinite geometric series where a1 = 65 and the common ratio is 1/6. Find the sum of this infinite series that will be the upper limit of this population. 78 28 11 32

OpenStudy (anonymous):

@hartnn

OpenStudy (campbell_st):

the formula for infinite sum is \[S_{\infty} = \frac{a_{1}}{1 - r}\] just substitute you're values of a1 and r.... to get the solution

OpenStudy (anonymous):

yes if you use the formula you get 65/1-(1/6) which is 65/(5/6) or (65*6)/5 which is 390/5=78

OpenStudy (anonymous):

Thank you @maida.beganovic

OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)^n − 1. 1,223 −1,023 −7,812 7,812

OpenStudy (anonymous):

@maida.beganovic

OpenStudy (anonymous):

Do I plug in all the numbers that are from 1-6 and add the answers together? @maida.beganovic

OpenStudy (anonymous):

I got -7812! is that the correct answer? @maida.beganovic

OpenStudy (anonymous):

yes it's correct but i think you should use the formula because you can only solve it this way for smaller numbers because there are only 6 members of this sequence that you had to add up :)

OpenStudy (anonymous):

Find the sum of the summation of 2 i minus 12, from i equals 7 to 16. 22 110 220 440

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

What did you get for the last one??